OpenStudy (anonymous):

how do you answer this question: Write the property of addition in following number sentence. 78+0=78

OpenStudy (mokeira):

Have you copied the question correctly

OpenStudy (opcode):

The property of addition in respect to the equation of $$ 78 + 0 = 78 $$ is the additive identity property. The additive identity property states that adding a number to 0 leaves it unchanged.

OpenStudy (anonymous):

hi mokeira.. yes this is what's in my sons assignment..

OpenStudy (anonymous):

hi opcode.. so how will i answer it, i have few more like 8+2=2+8

OpenStudy (mokeira):

basically, the propert of addition is something like a+b=b+a so i guess it will be 78+0=0+78

OpenStudy (opcode):

Give me a moment to write out all the properties of addition. :-)

OpenStudy (anonymous):

any idea what number sentence is?

OpenStudy (opcode):

\(\textbf{Commutative property}\): When two numbers are added, the sum is the same regardless of the order of the addends. For example: \(a + b = b + a\) \(\textbf{Associative Property\): When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example: \((2 + 3) + 4 = 2 + (3 + 4)\) \(\textbf{Additive Identity Property\): The sum of any number when added to zero is the original number. For example: \(a + 0 = a\). \(\textbf{Distributive property\): The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example: \(4 * (6 + 3) = 4*6 + 4*3\)

OpenStudy (mokeira):

In mathematics education, a number sentence is typically an equation or inequality expressed using numbers and common symbols.

OpenStudy (anonymous):

whooo.. to say think that this is just for grade one students... i need to go back to school... thanks so much for all your help.. :)

OpenStudy (opcode):

No problem, I am glad to help.

OpenStudy (mokeira):

:) u welcome @charmedkathy

OpenStudy (anonymous):

a+0=0+a